OpenStudy (anonymous):

In a multiple-choice test, each question has four options. Students will get 10 points for each correct answer; lose 4 points for each incorrect answer; and receive no points for unanswered questions. A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer? Show your calculations for expected value.

OpenStudy (anonymous):

each has probability 1/4 = .25 multiply and add. if you don't guess you get nothing. if you do guess you are right 1/4 of the time and gain 10 points, are wrong 3/4 of the time and lose 4 points so if you guess expected value is \[10\times \frac{1}{4}-4\times \frac{3}{4}=\frac{10-12}{4}=\frac{-2}{4}=-\frac{1}{2}\]

OpenStudy (anonymous):

on average each guess loses you half a point. don't guess

OpenStudy (anonymous):

hello.... master!

OpenStudy (amistre64):

howdy :)

OpenStudy (anonymous):

how'd i do here? hope it is clear

OpenStudy (anonymous):

i take you two know each other. LOL!!

OpenStudy (anonymous):

intimate friends, whoever he is

OpenStudy (anonymous):

lol! what

OpenStudy (amistre64):

...... know who? :)

OpenStudy (anonymous):

you two

OpenStudy (anonymous):

who knows? ;)

OpenStudy (amistre64):

lol.... i feel like this should be my dojo. Wax on, wax off...

OpenStudy (anonymous):

the waxer

OpenStudy (anonymous):

i meant my the greeting that you two gave to each other. keep up with the pace people.

OpenStudy (amistre64):

..if you can take this peeble from my hand; then... no,no,no... i wasnt ready!

OpenStudy (anonymous):

is it my fault site is so slow? ahh very good grasshopper...

OpenStudy (anonymous):

life is like ball of string...

OpenStudy (amistre64):

satellite, go.. bring me soup; but DONT touch it lol

OpenStudy (anonymous):

you two are a VERY interesting match.

OpenStudy (amistre64):

:) didja get your answer for the value question

OpenStudy (anonymous):

yeah its -1/2 right and it would be to her disadvantage to guess

OpenStudy (amistre64):

good :) value*probability + value*probability + .... = solution

OpenStudy (anonymous):

10×1/4 −4×3/4 =10−12/ 4 =−2/4 = -1/ 2